# Convergence of the extended Kalman filter with small and state-dependent noise

### v1

@@ Abstract
Nonlinear filtering problems are encountered in many applications, and one solution approach is the extended Kalman filter, which is not always convergent. Therefore, it is crucial to identify conditions under which the extended Kalman filter provides accurate approximations. This paper generalizes two significant results from picard1991efficiency on the efficiency of the continuous-time extended Kalman filter to a more general setting where the observation noise may be state-dependent but does not allow signal reconstruction from the quadratic variation of the observation process as in epidemic models. Firstly, we show that when the observation’s drift coefficient is strongly injective and the signal’s and observation’s drift become nearly linear for the diffusion scaling coefficient $\epsilon\to 0$, the estimation error is of order $\sqrt{\epsilon}$. Subsequently, we establish conditions under which the impact of the initial filtering error decays exponentially fast.
@@ 1 Introduction
This research investigates a nonlinear filtering problem in the presence of state-dependent noise. Many physical phenomena are best described by nonlinear sto-
chastic dynamics, in which the diffusion coefficient depends on the system’s state. These characteristics are essential for accurately modeling and understanding such systems. Therefore, filtering problems associated with these phenomena are nonstandard and necessitate alternative methods to the standard Kalman filtering approach. For instance, estimating the angular procession of a rotating spacecraft requires precise nonlinear filtering techniques to account for the state-dependent variations in noise. Similarly, the design of phase-locked loops, which are essential in communication systems to maintain signal synchronization, also requires these advanced filtering methods 1099828. These examples highlight the practical significance and wide-ranging applications of our study. The model that we will be interested in is the following.
$\begin{split}dY(t)&=f(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\sigma(t,Y(t),\pazocal{Z}(t))dW^{(1)}(t)+\sqrt{\varepsilon}g(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\\[2.15277pt] dZ(t)&=h(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\ell(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\end{split}$ (1)
with $(Y(0),Z(0))=(y_{0},z_{0})$, and where $Y(t)\in\mathbb{R}^{n}$ is our hidden signal and $Z(t)\in\mathbb{R}^{d}$ is our observation process at time $t\geq 0$. Additionally, $\pazocal{Z}(t)=\{Z_{s},s\leq t\}$ denotes the path of the observation process until the time point $t$. Finally, $\varepsilon$ is a small positive constant and $W^{(i)},~i=1,2$ are two independent standard Wiener processes in $\mathbb{R}^{d_{1}}$and $\mathbb{R}^{d_{2}}$ respectively (with $n,d,d_{1},d_{2}$ being positive integers).
This model is quite similar to the model in picard1991efficiency. However, we address a more general setting by allowing the diffusion coefficient of the observation process to be non-constant and state-dependent. For such a model, when the quadratic variation of the observation process is informative and can be computed in practice, it can be possible to reconstruct the signal under some injectivity conditions. This reconstruction is often not possible as in the context of epidemic models for which injectivity conditions are not fulfilled. It is therefore important to know whether the extended Kalman filter is efficient for such filtering problems.
@@
Many papers in the literature have been devoted to this type of filtering problem for which the diffusion coefficient of the observation is state-dependent. For instance, mclane1969optimal have proposed an optimal linear filter by restricting the filter structure to the linear class with respect to the observation. takeuchi1981least; takeuchi1985least have provided a more accurate approximation for the least-square optimal estimate as well as the convergence property of that estimate. More recently, joannides1995nonlinear; joannides1997nonlinear have considered nonlinear filtering with perfect observation and noninformative quadratic variation, which helps design robust and efficient algorithms in the case of filtering with small observation noise. Moreover, the case of state-dependent and small noise for Markovian diffusion has been investigated by picard1993estimation and he suggested a stochastic algorithm for the approximation of the optimal estimate.
On the other hand, several authors have considered the nonlinear filtering problem with small noise. Most of them are interested in finding efficient asymptotic approximate filters katzur1984asymptotic1; katzur1984asymptotic2. picard1991efficiency proved some convergence results on the extended Kalman filter for filtering problems with constant and small diffusion coefficient in the observation process dynamic.
@@
The main contributions of this paper are the generalization of two important results on error estimates of the extended Kalman filter from picard1991efficiency. First, we prove that the normalized the filtering error is of order $\sqrt{\varepsilon}$ for the more general model 1 under some assumptions. Second, we show that the filtering error is bounded over time, and the impact of the initial error on the filtering error vanishes exponentially fast.
@@
In Section 2, we introduce a motivating example from mathematical epidemiology, then Section 3 describes the derivation of the extended Kalman filter for (1), and presents some relevant definitions related to dynamical systems. Section 4 focuses on the estimation error of the extended Kalman filter, and we prove that under some assumptions, the filtering error is of order $\sqrt{\varepsilon}$. Finally, in Section 5, we show that the filtering error is bounded, and the initial error is forgotten exponentially fast when certain detectability conditions are met. The Appendix collects auxiliary results needed for the proofs of our theorems.
2 Motivating example: $SI^{\pm}S$ epidemic model
@@
In this section, we introduce a filtering system from mathematical epidemiology that is related to stochastic epidemic models with partial information and can be reformulated using appropriate time scaling as the filter system specified in (1). Let us consider the $SI^{\pm}S$ epidemic model in which a population is divided into three compartments according to the health state of individuals: susceptible ($S$), undetected infected ($I^{-}$), and detected infected ($I^{+}$). The associated flowchart is depicted in Figure 1.
Figure 1: Flowchart of the $SI^{\pm}$ compartmental model.
This model assumes that when a susceptible individual is infected, he or she is first presymptomatic or asymptomatic. This individual is accounted for in the $I^{-}$ compartment. After a test that occurs at a rate of $\alpha$, the undetected infected individuals become detected and move to the $I^{+}$ compartment. Here, people are in quarantine and can no longer infect new susceptible individuals. Both detected and undetected infected individuals recover from the infection and return to the susceptible compartment. The parameters $\beta$, $\rho^{-}$, and $\rho^{+}$ are the infection intensity, the recovery rates for the undetected and detected infected, respectively. We consider the population size $N$ to be constant, that is, $S+I^{-}+I^{+}=N$. A more detailed description of such models can be found in kamkumo2025estimating; kamkumo2025stochastic; ouabo2025phd, njiasse2025stoch; mbouandi2025phd. The stochastic dynamics of such a compartmental model can be derived using a continuous-time Markov chain approach as described in BrittonPardoux2019 and in kamkumo2025stochastic. For large population sizes $N$, diffusion approximations can be derived for the state vector $X$, which contains the relative subpopulation sizes. These are stochastic differential equations of the form
$dX(t)$ $=\overline{F}(X(t))dt~+~\dfrac{1}{\sqrt{N}}\overline{\sigma}(X(t))dB(t),~~X(0)=x_{0},$ (2)
@@
where coefficients $\overline{F}$, $\overline{\sigma}$ are non-linear functions and $B$ is a vector of independent standard Brownian motions. For instance, in the case of the above $SI^{\pm}S$ epidemic model, if we work with the reduced state $X=(X_{1},X_{2})^{\top}=(\overline{I}^{-},\overline{I}^{+})^{\top}=N^{-1}({I}^{-},{I}^{+})^{\top}$, Since we assume a constant population size, the proportion of susceptible individuals can be removed from the state, as it can be obtained from $\overline{S}=1-\overline{I}^{+}-\overline{I}^{-}$). The expressions of the functions $\overline{F}$ and $\overline{\sigma}$ read as follows, when the state $X$ is split into hidden component $Y=\overline{I}^{-}$ and observable component $Z=\overline{I}^{+}$:
$\overline{F}\left(\begin{array}[]{l}y\\ z\end{array}\right)$ $=\left(\begin{array}[]{l}\beta(1-y-z)y-(\alpha+\rho^{-})y\\[4.30554pt] \alpha y-\rho^{+}z\end{array}\right),$ (7)
@@
and $B=\left({B^{(1)}},\ldots,{B^{(4)}}\right)^{\top}$.
Let $\varepsilon={1}/{\sqrt{N}}$, $\tau={t}/{\varepsilon}=\sqrt{N}t$ and $U^{\varepsilon}:=\varepsilon^{-1}\left(X(\varepsilon\tau)-x_{0}\right)$. Using Equation (2) to derive the dynamics of the state $U^{\varepsilon}$ under the time scaling $t=\varepsilon\tau$, we get
$dU^{\varepsilon}(\tau)$ $=\overline{F}(x_{0}+\varepsilon U^{\varepsilon}(\tau))d\tau~+~\sqrt{\varepsilon}\overline{\sigma}(x_{0}+\varepsilon U^{\varepsilon}(\tau))dW(\tau),~~U^{\varepsilon}(0)=0,$ (14)
where $W(\tau)=\varepsilon^{-1/2}B(\varepsilon\tau)$. The scaling property of Brownian motion implies that the components of $W$ are independent standard Brownian motions. On a small time window with length depending on $\varepsilon$, functions $\overline{F}$ and $\overline{\sigma}$ can be approximated by linear functions since in the timescale $\tau$ their respective nonlinearities are less pronounced. The filtering system obtained for the components of $U^{\varepsilon}=(Y,Z)^{\top}$ with hidden signal $Y$ and observation $Z$ has exactly the structure of the system (1), with following functions coefficients:
$f(t,y,z)~=$ $~\beta(1-y-z)y-(\alpha+\rho^{-})y,\quad h(t,y,z)~=~\alpha y-\rho^{+}z,$ (15)
$\sigma(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\beta(1-y-z)y},&-\sqrt{\rho^{-}y}\end{array}\right),\quad g(t,y,z)~=~\left(\begin{array}[]{cc}-\sqrt{\alpha y},&0\end{array}\right),$ (18)
$\ell(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\alpha y},&-\sqrt{\rho^{+}z}\end{array}\right),$ (20)
@@ 3.1 Extended Kalman filter
The Kalman filtering procedure is well known and has been successfully applied in many domains. However, it requires the filtering problem to be linear. Some attempts have been made to apply the Kalman filter core idea to nonlinear models. One of the most prominent attempts is the so-called extended Kalman filter. This extended approach was first introduced by gelb1974applied, and an empirical justification was proposed by pardoux1991filtrage. This approach is described in many books, for instance bain2008fundamentals for more details. In the following computations and the rest of this work, functions will be of three variables: time $t$, signal $Y(t)$ and observation path history $\pazocal{Z}(t)$. However, for the sake of short notation, we will in some expressions omit dependence on the observation path history. Let us revisit the main step of the derivation of the extended Kalman filter for the nonlinear filtering problem (1). Let us assume that $\overline{Y}$ is the solution of the associated ordinary differential equation to our signal stochastic differential equation obtained by removing the diffusion term,
$\dfrac{d\overline{Y}(t)}{dt}=f(t,\overline{Y}(t)),~~~~\overline{Y}(0)=y_{0}.$ (21)
On a small interval of time $0,\Delta[$, with $\Delta>0$, we can perform the Taylor-like expansion of our nonlinear filtering system drift around the solution $\overline{Y}$, and replace $Y$ by $\overline{Y}$ in the diffusion coefficients. Hence, we obtain the linearized filtering system
@@
where $\nabla_{y}f$ and $\nabla_{y}h$ are the derivatives of $f$ and $h$ with respect to the signal argument, $Y$ respectively. Since the latter system is linear in $Y$, one can argue that in a small interval of time the equations satisfied by the original signal $Y$ and observation $Z$ are nearly linear. This setting is then suitable for the result in LiptserShiryaevVolII2001 on conditionally Gaussian process. By applying it, we compute from the linearized problem the approximations of the conditional mean $\widehat{Y}$ and the conditional variance $Q^{\varepsilon}$ that satisfy
@@
$+\int_{0}^{t}G(s)\big[dZ(s)-(h(s,\widehat{Y}(s))-\nabla_{y}h(s,\overline{Y}(s))\overline{Y}(s)+\nabla_{y}h(s,\overline{Y}(s))\widehat{Y}(s))ds\big],$ (22)
@@
$G(t)=\big[\varepsilon g\ell^{\top}(t,\overline{Y}(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,\overline{Y}(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,\overline{Y}(t)),$ (23)
and
@@
$+\nabla_{y}f(t,\overline{Y}(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,\overline{Y}(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,\overline{Y}(t)).$ (24)
By this method, it is possible to construct a mapping which relates any observable process of linearization such as $\overline{Y}$ to the corresponding conditional mean of the filter. Then, the extended Kalman filter by definition is the fixed point for this mapping, see bain2008fundamentals. Hence, we deduce that the extended Kalman filter for the nonlinear filtering problem (1) is given by the conditional mean $M$ that satisfies
$M(t)~=$ $M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (25)
@@
$G(t)=\big[\varepsilon g\ell^{\top}(t,M(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,M(t)),$ (26)
@@
$\dot{Q^{\varepsilon}}(t)=$ $~-\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]^{\top}(t,M(t))$ (27)
$~+\nabla_{y}f(t,M(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,M(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,M(t)).$ (28)
We can now define what will be called in the sequel, the extended Kalman filter for the filtering system (1). For the sake of simpler equations, we introduce the scaled covariance matrix $Q=\dfrac{1}{\varepsilon}Q^{\varepsilon}$.
Definition 3.1
Let $(M(t))_{t\geq 0}$ and $(Q^{\varepsilon}(t))_{t\geq 0}$ be two observable processes with values respectively in $\mathbb{R}^{n}$ and the set of positive-definite symmetric matrices of order $n$. A process which at any time $t$ is Gaussian with mean $M(t)$ and covariance matrix $Q^{\varepsilon}(t)=\varepsilon Q(t)$ will be said to be an extended Kalman filter for (1) if $M(t)$ is solution of
$M(t)=M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (29)

### v2

@@ Abstract
Nonlinear filtering problems are encountered in many applications, and one solution approach is the extended Kalman filter, which is not always convergent. Therefore, it is crucial to identify conditions under which the extended Kalman filter provides accurate approximations. This paper generalizes two significant results of Picard (1991) on the efficiency of the continuous-time extended Kalman filter for a filtering system with small noise, to a more general setting where the observation noise may be state-dependent but does not allow signal reconstruction from the quadratic variation of the observation process as for example in epidemic models. First, we show that if the drift of the signal process and the observation process becomes nearly linear when the parameter $\epsilon$, which scales the diffusion coefficients, approaches zero, and the drift coefficient of the observation process is strongly injective, then the estimation error is of the order of $\sqrt{\epsilon}$. We then establish conditions under which the impact of the initial filtering error decays exponentially fast.
Keywords Nonlinear filtering, small noise, state-dependent noise, extended Kalman filter, error estimate
2020 Mathematics Subject Classification 60G35, 62M05, 62M15, 93Ell
footnotetext: Corresponding author
1Brandenburg University of Technology Cottbus-Senftenberg, Institute of Mathematics,
P.O. Box 101344, 03013 Cottbus, Germany
Email: Ibrahim.MbouandiNjiasse@b-tu.de,   Florent.OuaboKamkumo@b-tu.de,  ralf.wunderlich@b-tu.de
@@ 1 Introduction
This research investigates a nonlinear filtering problem in the presence of state-dependent noise. Many physical phenomena can best be described by nonlinear stochastic differential equations (SDEs), in which not only the drift coefficient but also the diffusion coefficient depends on the state of the system. These characteristics are essential for accurately modeling and understanding such systems. Filtering problems associated with these phenomena are nonstandard and necessitate alternative methods to the standard Kalman filtering approach. For instance, estimating the angular procession of a rotating spacecraft requires precise nonlinear filtering techniques to account for the state-dependent variations in noise. Similarly, the design of phase-locked loops, which are essential in communication systems to maintain signal synchronization, also requires these advanced filtering methods [10]. Another application from mathematical epidemiology is considered below in Section 2. These examples highlight the practical significance and wide-ranging applications of our study.
The general setting of the stochastic filtering problem can be expressed in terms of the following coupled system of SDEs on $0,\infty)$
$\begin{split}dY(t)&=f(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\sigma(t,Y(t),\pazocal{Z}(t))dW^{(1)}(t)+\sqrt{\varepsilon}g(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\\[2.15277pt] dZ(t)&=h(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\ell(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\end{split}$ (1.1)
with initial values $(Y(0),Z(0))=(y_{0},z_{0})$. Here, $Y(t)\in\mathbb{R}^{n}$ denotes the hidden signal process and $Z(t)\in\mathbb{R}^{d}$ the observation process at time $t\geq 0$. Further, $\pazocal{Z}(t)=\{Z_{s},s\leq t\}$ denotes the path of the observation process up to time $t$. Finally, $\varepsilon>0$ is a small constant that scales the diffusion coefficients, and $W^{(i)},~i=1,2$, are two independent standard Brownian motions in $\mathbb{R}^{d_{1}}$and $\mathbb{R}^{d_{2}}$ respectively (with $n,d,d_{1},d_{2}$ being positive integers).
This model is quite similar to the model in [22]. However, we address a more general setting by allowing the diffusion coefficient of the observation process to be non-constant and state-dependent. For such a model, when the quadratic variation of the observation process is informative and can be computed in practice, it can be possible to reconstruct the signal under some injectivity conditions. However, this reconstruction is often not possible, as in the case of epidemic models, see Section 2, for which the injectivity conditions are not fulfilled. It therefore remains an open question, which we investigate in this paper, whether the results of [22] on the efficiency of the extended Kalman filter can be transferred to the generalized setting in (1.1).
@@
Filtering problems with state-dependent diffusion coefficients in the SDE of the observation process are for instance studied in [9]. The author proposed an optimal linear filter by restricting the filter structure to the linear class with respect to the observation. Further references are [7, 8], which provided a more accurate approximation for the least-square optimal estimate as well as the convergence property of that estimate. More recently, [5, 6] have considered nonlinear filtering with perfect observation and noninformative quadratic variation, which helps design robust and efficient algorithms in the case of filtering with small observation noise. As mentioned above, the case of state-dependent and small noise for Markovian diffusion has been investigated by [23] and he suggested a stochastic algorithm for the approximation of the optimal estimate.
On the other hand, several authors have considered the nonlinear filtering problem with small noise. Most of them are interested in finding efficient asymptotic approximate filters, see [3, 4]. [22] proved some convergence results on the extended Kalman filter for filtering problems with constant and small diffusion coefficient in the observation process dynamic. More recently, [11] considered such a filtering problem with small observation noise, and designed an adaptive extended Kalman filter algorithm that estimates both an unknown parameter and the unobservable state.
@@
The main contributions of this paper are the generalization of two important results on error estimates of the extended Kalman filter from [22]. Under certain suitable assumptions, we first show that the normalized filter error for the generalized model (1.1) is of the order of $\sqrt{\varepsilon}$. Second, we show that the filtering error is bounded over time, and the impact of the initial error on the filtering error vanishes exponentially fast.
@@
In Section 2, we introduce a motivating example from mathematical epidemiology, then Section 3 describes the derivation of the extended Kalman filter for (1.1), and presents some relevant definitions related to dynamical systems. Section 4 focuses on the estimation error of the extended Kalman filter, and we prove that under some assumptions, the filtering error is of order $\sqrt{\varepsilon}$. Finally, in Section 5, we show that the filtering error is bounded, and the initial error is forgotten exponentially fast when certain detectability conditions are met. The Appendix collects auxiliary results needed for the proofs of our theorems.
2 Motivating example from mathematical epidemiology
Figure 1: Flowchart of the $SI^{\pm}S$ compartmental model.
@@ 2 Motivating example from mathematical epidemiology
In this section, we introduce a filtering system from mathematical epidemiology that is related to stochastic epidemic models with partial information and can be approximated by the filter system specified in (1.1). Let us consider the $SI^{\pm}S$ epidemic, modeling the spread of an infectious disease in which a population is divided into three compartments according to the health state of individuals: susceptible ($S$), undetected infected ($I^{-}$), and detected infected ($I^{+}$). The associated flowchart is depicted in Figure 1. Individuals that are susceptible to the disease form the compartment $S$. By contact to infectious individuals they become infected and transition with rate $\beta>0$ to the compartment $I^{-}$ of undetected infected individuals. Here, presymptomatic or asymptomatic individuals and those who are not aware about their infection are collected. All of them are considered infectious and are not in quarantine. The health authorities test the population for the infectious disease with a test rate $\alpha>0$. People from group $I^{-}$ test positive and move to group $I^{+}$, which consists of confirmed or detected infected individuals who are in quarantine and can no longer infect susceptible individuals. Both detected and undetected infected individuals recover from the infection with recovery rates $\rho^{+},\rho^{-}>0$, and return to the susceptible compartment $S$. A more detailed description of such models can be found in [20, 19, 18], [16, 17].
We denote the number of individuals in the three compartments at time $t\geq 0$ by $S(t),I^{-}(t),I^{+}(t)$. The total population size is assumed to be constant and denoted by $N$. Then for all $t\geq 0$ it holds $S(t)+I^{-}(t)+I^{+}(t)=N$. Since the number of confirmed infected individuals in compartment $I^{+}$ is known to health authorities due to monitoring through testing, $I^{+}(t)$ can be treated as an observable variable. However, the size of the compartments $S$ and $I^{-}$ is not directly observable, as the infection status of their individuals is unknown. Therefore, $S(t)$ and $I^{-}(t)$ must be treated as hidden or unobservable variables that can only be estimated based on observations of $I^{-}$.
The stochastic dynamics of such a compartmental model can be derived using a continuous-time Markov chain approach as described in [2] and in [19]. For large population sizes $N$, diffusion approximations can be derived for the state vector $X$, which contains the relative subpopulation sizes. These are stochastic differential equations of the form
$dX(t)$ $=\overline{F}(X(t))dt~+~\dfrac{1}{\sqrt{N}}\overline{\sigma}(X(t))dW(t),~~X(0)=x_{0},$ (2.1)
where coefficients $\overline{F}$, $\overline{\sigma}$ are nonlinear functions and $W$ is a vector of independent standard Brownian motions.
@@
In the case of the above $SI^{\pm}S$ epidemic model, we can work with the reduced state $X=(X_{1},X_{2})^{\top}=(\overline{I}^{-},\overline{I}^{+})^{\top}=N^{-1}({I}^{-},{I}^{+})^{\top}$. Since we assume a constant population size, the proportion of susceptible individuals $\overline{S}=S/N$ can be removed from the state, as it can be derived from $\overline{S}=1-\overline{I}^{+}-\overline{I}^{-}$. If the state $X$ is divided into the hidden component $Y=\overline{I}^{-}$ and the observable component $Z=\overline{I}^{+}$, and the notation Y(t)=y,Z(t)=zY(t)=y,~Z(t)=z and $\overline{S}(t)=s$ is used, the expressions of the functions $\overline{F}$ and $\overline{\sigma}$ are as follows, see [17, 19]:
@@
and $W=\left({W_{1}},\ldots,{W_{4}}\right)^{\top}$ with independent standard Brownian motions $W_{1},\ldots,W_{4}$.
The expression $sy$ or, after replacing the variable $s$ with 1yz1-y-z, the term (1yz)y(1-y-z)y leads to a quadratic nonlinearity in this dynamic, in particular to a nonlinear dependence of the drift coefficient on the signal $y$. This makes the filter problem nonlinear, and in this form it does not meet the technical conditions required for the subsequent analysis. Therefore, we first divide the time interval $[0,T]$ into $N_{t}\in\mathbb{N}$ subintervals of length $\Delta t=T/N_{t}$ and grid points $t_{i}=i\Delta t,i=0,\ldots,N_{t}$. Second, we are going to approximate the dynamics of the state $X=(Y,Z)^{\top}$ on the $N_{t}$ small time intervals $t_{i},t_{i+1}),~i=0,\ldots,N_{t}-1$, by freezing the variable $s$ to its value at the left endpoint of the interval $s_{i}=\overline{S}(t_{i})$. This is motivated by the fact that in epidemic models, the proportion of susceptible individuals is usually much larger than the proportions in the other compartments. Therefore, compared to the proportions in the rather “small” compartments $I^{-}$ and $I^{+}$, it varies much more slowly and can be approximated well by a constant, in contrast to the latter variables, which are of greatest interest in epidemic models.
This yields a filtering system on the interval $t_{i},t_{i+1}),~i=0,\ldots,N_{t}-1$ that has linear drift coefficients and nonlinear diffusion coefficients. Moreover, this filtering system has exactly the structure of the system (1.1), with $\varepsilon={1}/{N}$ and the following coefficient functions:
$f(t,y,z)~=$ $~\beta s_{i}y-(\alpha+\rho^{-})y,$ $h(t,y,z)$ $=~\alpha y-\rho^{+}z,$ (2.11)
$\sigma(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\beta s_{i}y},&-\sqrt{\rho^{-}y}\end{array}\right),$ $g(t,y,z)$ $=~\left(\begin{array}[]{cc}-\sqrt{\alpha y},&0\end{array}\right),$ (2.14)
$\ell(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\alpha y},&-\sqrt{\rho^{+}z}\end{array}\right),$ (2.16)
@@ 3.1 Extended Kalman filter
The Kalman filtering procedure is well known and has been successfully applied in many domains. In its standard form, however, it is necessary that, in the SDEs of the filtering system, the drift coefficients be linear in the signal and the diffusion coefficients be independent of the signal. Some attempts have been made to apply the core idea of the Kalman filter to nonlinear models. One of the most prominent attempts is the so-called extended Kalman filter.
Let us fix a filtered probability spaces, denoted by $\left(\Omega,\pazocal{F},\mathbb{F},\mathbb{P}\right)$. The filtration $\mathbb{F}=(\pazocal{F}_{t})_{t\geq 0}$ with $\pazocal{F}_{s}\subset\pazocal{F}_{t}\subset\pazocal{F}$ for $0\leq s<t$, is such that all stochastic processes appearing in throughout this work are adapted with respect to it. The extended Kalman filter aims to approximate the mean-square optimal estimate of the signal given the observations which is known to be the projection of the signal process onto the the filtration $\mathbb{F}^{Z}\subset\mathbb{F}$ generated by the observation process $(Z(t))_{t\geq 0}$, i.e., $\mathbb{F}^{Z}=(\pazocal{F}^{Z}_{t})_{t\geq 0}$ with the $\sigma$-algebras $\pazocal{F}^{Z}_{t}=\sigma\{Z(s),s\leq t\}$. This estimate is given by the the conditional mean $\widehat{M}(t)=\mathbb{E}[Y(t)|\pazocal{F}^{Z}_{t}]$. The associated estimation error is measured by the conditional covariance matrix $\mathrm{Var}[Y(t)|\pazocal{F}^{Z}_{t}]=\mathbb{E}[(Y(t)-\widehat{M}(t))(Y(t)-\widehat{M}(t))^{\top}|\pazocal{F}^{Z}_{t}].$ This extended approach was first introduced by [12], and an empirical justification was proposed by [21]. This approach is described in many books, for instance [1] for more details.
Remark 3.1.
In the following, many functions will have three variables: time $t$, signal $Y(t)$, and observation path $\pazocal{Z}(t)$, as already seen in the coefficients of the SDEs in (1.1). For the sake of brevity, however, we will omit the dependence on the observation path $\pazocal{Z}(t)$ in some expressions, but keep this dependence in mind.
Let us revisit the main step of the derivation of the extended Kalman filter for the nonlinear filtering problem (1.1). We denote by $\overline{Y}$ the solution of the ordinary differential equation associated with the SDE for the signal process $Y$ in (1.1), and obtained by removing the diffusion term, i.e.
$\dfrac{d\overline{Y}(t)}{dt}=f(t,\overline{Y}(t)),~~~~\overline{Y}(0)=y_{0}.$ (3.1)
In a small interval of time $0,\Delta)$, with $\Delta>0$, we can perform the Taylor-like expansion of the nonlinear drift coefficients of the filtering system (1.1) around the solution $\overline{Y}$, and replace in the diffusion coefficients $Y$ with $\overline{Y}$. Then, we obtain the following linearized filtering system
@@
where $\nabla_{y}f$ and $\nabla_{y}h$ denote the gradients of $f$ and $h$ with respect to the signal argument $Y$, respectively. In a small interval of time, the drift coefficients of the SDEs in the original system can be assumed to be nearly linear in the signal $Y$, such that the above linearized SDEs provide a good approximation. This setting is then suitable to use the results from [13, Chapter 12] on conditionally Gaussian process. By applying it, we obtain approximations of the conditional mean $\widehat{Y}$ and the conditional variance $Q^{\varepsilon}$ by the corresponding filter processes for the linearized problem that satisfy, see [13, Theorem 12.66]
@@
$+\int_{0}^{t}G(s)\big[dZ(s)-(h(s,\widehat{Y}(s))-\nabla_{y}h(s,\overline{Y}(s))\overline{Y}(s)+\nabla_{y}h(s,\overline{Y}(s))\widehat{Y}(s))ds\big],$ (3.2)
@@
$G(t)=\big[\varepsilon g\ell^{\top}(t,\overline{Y}(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,\overline{Y}(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,\overline{Y}(t)),$ (3.3)
and $Q^{\varepsilon}$ satisfies the Riccati differential equation
@@
$+\nabla_{y}f(t,\overline{Y}(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,\overline{Y}(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,\overline{Y}(t)).$ (3.4)
Note that this is an ordinary differential equation with random coefficients, as these depend on the observation path $\pazocal{Z}(t)$, which has been omitted in the notation. By this method, it is possible to construct a mapping which relates any observable process of linearization such as $\overline{Y}$ to the corresponding conditional mean of the filter. Then, the extended Kalman filter by definition is the fixed point for this mapping, see [1]. Hence, we deduce that the extended Kalman filter for the nonlinear filtering problem (1.1) is given by the conditional mean $M$ that satisfies
$M(t)~=$ $M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (3.5)
@@
$G(t)=\big[\varepsilon g\ell^{\top}(t,M(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,M(t)),$ (3.6)
@@
$\dot{Q^{\varepsilon}}(t)=$ $~-\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]^{\top}(t,M(t))$ (3.7)
$~+\nabla_{y}f(t,M(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,M(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,M(t)).$ (3.8)
In summary, we now formally define what will be referred to below as the extended Kalman filter for the filter system (1.1). For the sake of simpler equations, we introduce the scaled covariance matrix $Q=\dfrac{1}{\varepsilon}Q^{\varepsilon}$.
Definition 3.2.
Let $(M(t))_{t\geq 0}$ and $(Q^{\varepsilon}(t))_{t\geq 0}$ be two observable processes with values respectively in $\mathbb{R}^{n}$ and the set of positive-definite symmetric matrices of order $n$. A process which at any time $t$ is Gaussian with mean $M(t)$ and covariance matrix $Q^{\varepsilon}(t)=\varepsilon Q(t)$ will be said to be an extended Kalman filter for (1.1) if $M(t)$ is solution of
$M(t)=M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (3.9)